OpenStudy (anonymous):

Solve the word problem. The formula s = gives the side length s of a regular pentagon with perimeter P. What is the side length of a regular pentagon with perimeter 7.5 feet? A. 1.25 ft B. 1.5 ft C. 2.5 ft D. 12.5 ft

OpenStudy (anonymous):

help or amswer

OpenStudy (jdoe0001):

The formula s = gives the side length ^ that looks pretty blank there

OpenStudy (anonymous):

b

OpenStudy (jdoe0001):

anyway \(\bf \textit{side length}=\cfrac{perimeter}{\textit{number of sides}} \\ \quad \\ \begin{array}{ccllll} penta&gon\\ \uparrow \\ \textit{5 sides} \end{array}\implies \textit{side length}=\cfrac{7.5}{5}\)

OpenStudy (anonymous):

soooo

OpenStudy (anonymous):

im not good at math ok

OpenStudy (jdoe0001):

so that's how you'd use the formula a pentagon has 5 sides perimeter is 7.5 there you have it :)

OpenStudy (anonymous):

7.5 is nopt an answer

OpenStudy (jdoe0001):

\(\bf \textit{side length}=\cfrac{perimeter}{\textit{number of sides}} \\ \quad \\ \begin{array}{ccllll} penta&gon\\ \uparrow \\ \textit{5 sides} \end{array}\implies \textit{side length}=\Large \cfrac{7.5}{5}\)

OpenStudy (anonymous):

B

OpenStudy (jdoe0001):

yeap 7.5/5 = 1.5

OpenStudy (anonymous):

thank you